# Probing Dark Matter Substructures with Free-Form Modelling: A Case Study of the ‘Jackpot’ Strong Lens

### v1

@@ 2.3 Simultaneous Inversion of Source and Potential Perturbations
In Section 2.2, we fix the source light to the result given by the macro model, fitting only for $\boldsymbol{\delta\psi}$ using equation 6. However, this process does not account for the covariance between the source light and $\boldsymbol{\delta\psi}$. It has been shown that the source light model can absorb the image residual left by the imperfect lens mass model (e.g. Vernardos & Koopmans, 2022; Nightingale et al., 2024). Therefore, a better approach is to fit the source light and $\boldsymbol{\delta\psi}$ simultaneously, allowing them to compete to reconstruct the observed lensing image. As demonstrated by Koopmans (2005); Vegetti & Koopmans (2009a), this can be achieved with simple manipulations on equation 6. Suppose we use a pixelized source model for the macro model, then the image residual left by the macro model, which appears on the left-hand side of equation 6, is equal to
@@
Each element $d_{j}$ in the vector $\boldsymbol{d}$ represents the brightness value of the lensing image at each pixel. Each element $s_{i}$ in the vector $\boldsymbol{s}$ represents the brightness value at each source pixel. $\mathbf{L_{s}}(\boldsymbol{\psi}_{p})$ is the lens mapping matrix with dimensions $[N_{d},N_{s}]$, which, when acting on $\boldsymbol{s}$, generates the lensed source image. The value of $\mathbf{L_{s}}(\boldsymbol{\psi}_{p})$ depends on the macro model’s lens mass model (denoted by its lensing potential $\boldsymbol{\psi}_{p}$) and the interpolation scheme of the pixelized source model. For a more thorough illustration of how to construct this matrix $\mathbf{L_{s}}(\boldsymbol{\psi}_{p})$, refer to Appendix B of Treu & Koopmans (2004). On the right-hand side of equation 6, the construction of the matrix $\mathbf{L_{\delta\psi}}$ requires knowledge of the source brightness distribution $S$, which can be obtained by interpolating the pixelized source reconstruction (denoted by $\boldsymbol{s}_{p}$), given by the macro model. Reorganizing equation 6, we have
@@
It is immediately clear that the matrix equation 21 is linear; therefore, all the mathematical frameworks shown in Section 2.2 to solve Equation 6 can be applied to Equation 21. All we need to do is to replace $\delta\boldsymbol{d}$ with $\boldsymbol{d}$, $\mathbf{L_{\delta\psi}}$ with $\mathbf{L}$, and $\boldsymbol{\delta\psi}$ with $\boldsymbol{r}$ (see also Vernardos & Koopmans, 2022). One issue worth noting is that the regularisation prior $\{\boldsymbol{g}_{r},\boldsymbol{\xi}_{r}\}$ for the vector $\boldsymbol{r}$ can be split into two independent parts: the source light regularisation $\{\boldsymbol{g}_{s},\boldsymbol{\xi}_{s}\}$ and the lensing potential perturbation regularisation $\{\boldsymbol{g}_{\delta\psi},\boldsymbol{\xi}_{\delta\psi}\}$. The regularisation functions for both $\boldsymbol{s}$ and $\delta\boldsymbol{\psi}$ also take a quadratic form, expressed as
@@
We note that solving the unknown perturbations in the lensing potential, whether based on equation 6 or 21, requires a non-zero source gradient (embedded in the $\boldsymbol{D_{s}}$ matrix). This implies that the potential correction method is, in principle, only applicable in the annular region where lensed emission is present. In practice, it is possible to use a larger mask that includes more regions without significant lensed emission and to let regularisation forcibly remove ill-posed solutions. However, as discussed in Section 5.1, this approach tends to be sub-optimal.
@@ 2.4.2 Configurations for Lensing Potential Perturbations
Solving the unknown lensing potential perturbation with Equation 6 or 21 requires knowledge of non-vanishing source gradient information, meaning the potential correction method only works in an annular region that tightly encloses the lensed source images. The pixelization grid for the free-form model of lensing potential perturbation is thus defined within this annular region, using a regular grid with a pixel size $N_{\delta\psi}$ times larger than that of the native lensing image [^4] This functionality is supported by the RegularDpsiMesh module in our code, where the factor parameter corresponds to $N_{\delta\psi}$., where $N_{\delta\psi}=1,2,\ldots$. The default $N_{\delta\psi}$ value used in this work is 2.
The Matérn regularisation (Stein, 1999; Rasmussen & Williams, 2006; Matern, 2013) is employed to control the smoothness of the pixelized lensing potential perturbation. For any two pixels of the lensing potential perturbation, denoted by their coordinates $(x_{m},y_{m})$ and $(x_{n},y_{n})$, the correlation between them can be described by
$C_{ij}(\rho,\nu)=\sigma^{2}\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\sqrt{2\nu}\frac{d_{ij}}{\rho}\right)^{\nu}K_{\nu}\left(\sqrt{2\nu}\frac{d_{ij}}{\rho}\right),$ (28)
where $\Gamma$ is the gamma function, and $K_{\nu}$ is the modified Bessel function of the second kind. $d_{ij}=\sqrt{(x_{m}-x_{n})^{2}+(y_{m}-y_{n})^{2}}$ represents the distance between these two pixels. The parameter $\rho$ defines the characteristic smoothing scale, while $\nu$ governs the order of smoothing, ensuring that the pixelized model reconstruction is $\left\lceil\nu\right\rceil-1$ times differentiable. The Exponential and Gaussian kernels used in Vernardos & Koopmans (2022) are special cases of the Matérn kernel, with $\nu=0.5$ and $\nu=+\infty$, respectively. The Hessian ($\mathbf{R}_{\delta\psi}$) of the regularisation function, using the Matérn scheme, is defined as follows:
@@ 3 Overview of Strong Lensing Data
Gaussian Random Field Normalization ($\log[A_{\mathrm{GRF}}]$) $-6.0$ $\left(-4.35^{0.03}_{-0.03}\right)$
@@ 3.1 Mock Lensing Data
We generate a set of mock lensing images that incorporate various perturbation signals in the lensing potential to evaluate the effectiveness of the potential correction code developed in this study. To control for variables, we ensure that the mass of the main lens galaxy and the source light remain consistent across all mock datasets, allowing only the perturbation signals in the lensing potential to vary. The mass distribution of the main lens galaxy is described by the elliptical power-law model (Tessore & Metcalf, 2015), whose dimensionless surface mass density (i.e. convergence) profile is expressed as
@@
Here, $31<\gamma<$ 3 represents the slope of the deprojected volume density, $\theta_{\mathrm{E}}$ denotes the Einstein radius, and $r$ signifies the elliptical radial distance from the centre of the lens galaxy. The elliptical radial distance is defined as $r(x,y)=\sqrt{x^{2}q+\frac{y^{2}}{q}}$, where $q$ is the minor-to-major axis ratio, and $\phi$ is the position angle measured counterclockwise from the positive x-axis. The parameters $q$ and $\phi$ are reparametrized into two ellipticity components, $\epsilon_{1}$ and $\epsilon_{2}$, defined as
@@
Four popular types of lensing potential perturbation signals, frequently used to evaluate the performance of free-form models (e.g. Vernardos & Koopmans, 2022; Galan et al., 2022; Biggio et al., 2023), are employed to construct mock datasets and test our potential correction code. These include: a spherical Navarro-Frenk-White (NFW) profile representing subhalos within the lens galaxy (Navarro et al., 1996), an external shear field caused by neighbouring galaxies or the cosmic large-scale structure (Keeton et al., 1997), a multipole perturbation induced by galaxy interactions such as mergers (Hao et al., 2006), and a Gaussian Random Field (GRF), which is expected to capture the collective lensing effects of numerous line-of-sight halos (Chatterjee & Koopmans, 2018; Bayer et al., 2023). The NFW model has a volume mass density profile of
@@
where $a_{m}$ represents the normalization factor, and $m$ and $\phi_{m}$ denote the multipole order and orientation, respectively (c.f. Congdon & Keeton, 2005; Taylor et al., 2017; Etherington et al., 2024; Paugnat & Gilman, 2025). The characteristics of a GRF are defined by its isotropic power spectrum, which follows a power-law relation:
@@ 3.2 Lensing Data for SLACS0946+1006
SLACS0946+1006 is a compound lens system with a rare triple source-plane configuration, characterized by a lens redshift of $z_{l}=0.222$ and source redshifts of $z_{s1}=0.609$, $z_{s2}=2.035$, and $z_{s3}=5.975$ (Smith & Collett, 2021). This system was initially pre-selected as a lens candidate from early-type galaxies in the SDSS survey, which exhibited abnormal emission lines potentially originating from high-redshift background sources (Bolton et al., 2006). The ACS camera on board the HST was subsequently employed to confirm the strong lensing nature of this system. The SLACS0946+1006 data analysed in this work were obtained using the ACS F814W band and then drizzled to $0.05\arcsec\text{pixel}^{-1}$, following the reduction procedures described in Bolton et al. (2008). The PSF was determined using the TinyTim software. For simplicity, we masked the outer Einstein ring, using only the innermost Einstein ring for potential correction. The lens light was pre-subtracted using the model results from He et al. (2024), which employs the Multipole Gaussian Expansion (MGE) model to represent the lens light; combined with a pixelized source model, the lens and source light are fit simultaneously to account for the covariance between the two. Subtracting the lens light using MGE typically outperforms multiple Sersic (Sérsic, 1963) fits in terms of image residuals. Figure 2 shows the lens-light-subtracted image of SLACS0946+1006.
@@ 4.1 Inverting Only the Potential Perturbation
Figure 3: Reconstruction of the lensing potential perturbation induced by a $10^{10}M_{\odot}$ NFW subhalo using the Matérn regularisation scheme, assuming that the true source and the image residuals induced by the perturber are perfectly known. Top-left: Image residual induced by the $10^{10}M_{\odot}$ NFW subhalo. Top-middle: Image residual reconstructed using the potential correction model. Top-right: Difference between the top-left and top-middle panels, normalised by the noise map. Bottom-left: Lensing potential map of the input perturber. Bottom-middle: Lensing potential map of the model perturber, reconstructed using the potential correction model. Bottom-right: Difference between the bottom-left and bottom-middle panels.
@@
Figure 5: Comparison of reconstructed lensing potential perturbations induced by an NFW subhalo (left column), external shear (middle-left column), $m_{4}$ multipole (middle-right column), and a Gaussian random field (GRF) (right column) with the ground truth. The first row presents the input lensing potential maps of perturbers. The second row shows the perturbative lensing potential maps derived from the potential correction model. The third row presents a quantitative analysis: for the NFW subhalo, external shear, and $m_{4}$ multipole, parametric models are fitted to the lensing potential map in the second row, and their best-fit parameter values are summarised in Table 3; for the GRF, the power spectra of the input (blue lines) and reconstructed (red lines) lensing potential maps are compared. Solid lines indicate power spectra calculated within the unmasked modelling region, while dashed lines represent power-law model fits.
@@
Parameter Names $\delta\boldsymbol{\psi}$-Only $(\delta\boldsymbol{\psi},\boldsymbol{s})^{\mathrm{Truth}}$ $(\delta\boldsymbol{\psi},\boldsymbol{s})^{\mathrm{Macro}}$ Ground Truths
@@
Mass ($\log_{10}[M_{200}/\mathrm{M}_{\odot}]$) $9.9665^{0.7155}_{-0.6685}$ $9.9151^{0.7237}_{-0.6488}$ $9.5192^{0.8085}_{-0.3837}$ 10.0
@@
Table 3: Quantitative comparison of reconstructed and input lensing potential perturbations. For subhalo, shear, and multipole models, parametric models are fitted to the reconstructed perturbations to assess agreement with the input. For the Gaussian random field, the power spectrum of the perturbations is compared. The labels $\delta\boldsymbol{\psi}$-Only, $(\delta\boldsymbol{\psi},\boldsymbol{s})^{\mathrm{Truth}}$, and $(\delta\boldsymbol{\psi},\boldsymbol{s})^{\mathrm{Macro}}$ denote inversions of only the lensing potential perturbations, inversions of both source and lensing potential perturbations starting from the exact true main lens mass and source light distributions, and inversions of both source and lensing potential perturbations starting from the macro model fit, respectively. Superscripts and subscripts represent the 68% credible interval.
@@
To better visualise the structure of a localised subhalo, the convergence map is more informative than the lensing potential map. The top-left and top-right panels of Figure 4 compare the convergence maps of the input perturber and the one derived by the potential correction model, respectively. Our potential correction method with Matérn regularisation successfully reconstructs the mass distribution of the input NFW subhalo qualitatively. In contrast, using Exponential or Gaussian regularisation, as proposed in Vernardos & Koopmans (2022), leads to under- or over-smoothed solutions that fail to recover the input NFW subhalo, as shown in the bottom-left and bottom-right panels of Figure 4. This demonstrates the superiority of Matern regularisation in reconstructing localised subhalo perturbations.
To enable a more quantitative assessment of the potential correction method, we fit a parametric model to the perturbative lensing potential map derived by the potential correction method for mock lenses with an NFW subhalo [^5] The NFW subhalo is a localised perturber. For this parametric model fitting, we only use the potential correction results in the central region (indicated by the black circle in Figure 4), as the outskirts are likely degraded by noise., external shear, and an $m_{4}$ multipole perturber, and evaluate whether the fitting results agree with the true input values. For the Gaussian random field, we compare the power spectrum of the model-derived lensing potential perturbations with the true input values. As shown in Figure 5, the lensing potential perturbations derived by the potential correction method (second row) are consistent with the true input values (first row). A quantitative diagnostic (bottom row), conducted either via parametric model fitting or power spectrum comparison, demonstrates that our potential correction method can accurately recover the input perturbations. The specific parametric values from this diagnostic are summarised [^6] It should be noted that, although the power spectrum properties of the input Gaussian random field are recovered reasonably well using the potential correction method, the model and the input true values do not perfectly align within the statistical uncertainty, a phenomenon also observed in (Vernardos & Koopmans, 2022). in Table 3.

### v2

@@ 2.3 Simultaneous Inversion of Source and Potential Perturbations
In Section 2.2, we fix the source light to the result given by the macro model, fitting only for $\boldsymbol{\delta\psi}$ using equation 6. However, this process does not account for the covariance between the source light and $\boldsymbol{\delta\psi}$. It has been shown that the source light model can absorb the image residual left by the imperfect lens mass model (e.g. Vernardos and Koopmans, 2022; Nightingale et al., 2024). Therefore, a better approach is to fit the source light and $\boldsymbol{\delta\psi}$ simultaneously, allowing them to compete to reconstruct the observed lensing image. As demonstrated by Koopmans (2005); Vegetti and Koopmans (2009a), this can be achieved with simple manipulations on equation 6. Suppose we use a pixelized source model for the macro model, then the image residual left by the macro model, which appears on the left-hand side of equation 6, is equal to
@@
Each element $d_{j}$ in the vector $\boldsymbol{d}$ represents the brightness value of the lensing image at each pixel. Each element $s_{i}$ in the vector $\boldsymbol{s}$ represents the brightness value at each source pixel. $\mathbf{L_{s}}(\boldsymbol{\psi}_{p})$ is the lens mapping matrix with dimensions $N_{d},N_{s}]$, which, when acting on $\boldsymbol{s}$, generates the lensed source image. The value of $\mathbf{L_{s}}(\boldsymbol{\psi}_{p})$ depends on the macro model’s lens mass model (denoted by its lensing potential $\boldsymbol{\psi}_{p}$) and the interpolation scheme of the pixelized source model. For a more thorough illustration of how to construct this matrix $\mathbf{L_{s}}(\boldsymbol{\psi}_{p})$, refer to Appendix B of Treu and Koopmans (2004). On the right-hand side of equation 6, the construction of the matrix $\mathbf{L_{\delta\psi}}$ requires knowledge of the source brightness distribution $S$, which can be obtained by interpolating the pixelized source reconstruction (denoted by $\boldsymbol{s}_{p}$), given by the macro model. Reorganizing equation 6, we have
@@
It is immediately clear that the matrix equation 12 is linear; therefore, all the mathematical frameworks shown in Section 2.2 to solve Equation 6 can be applied to Equation 12. All we need to do is to replace $\delta\boldsymbol{d}$ with $\boldsymbol{d}$, $\mathbf{L_{\delta\psi}}$ with $\mathbf{L}$, and $\boldsymbol{\delta\psi}$ with $\boldsymbol{r}$ (see also Vernardos and Koopmans, 2022). One issue worth noting is that the regularisation prior $\{\boldsymbol{g}_{r},\boldsymbol{\xi}_{r}\}$ for the vector $\boldsymbol{r}$ can be split into two independent parts: the source light regularisation $\{\boldsymbol{g}_{s},\boldsymbol{\xi}_{s}\}$ and the lensing potential perturbation regularisation $\{\boldsymbol{g}_{\delta\psi},\boldsymbol{\xi}_{\delta\psi}\}$. The regularisation functions for both $\boldsymbol{s}$ and $\delta\boldsymbol{\psi}$ also take a quadratic form, expressed as
@@
We note that solving the unknown perturbations in the lensing potential, whether based on equation 6 or 12, requires a non-zero source gradient (embedded in the $\boldsymbol{D_{s}}$ matrix). This implies that the potential correction method is, in principle, only applicable in the annular region where lensed emission is present. In practice, it is possible to use a larger mask that includes more regions without significant lensed emission and to let regularisation forcibly remove ill-posed solutions. However, as discussed in Section 5.1, this approach tends to be sub-optimal.
@@ 2.4.2 Configurations for Lensing Potential Perturbations
Solving the unknown lensing potential perturbation with Equation 6 or 12 requires knowledge of non-vanishing source gradient information, meaning the potential correction method only works in an annular region that tightly encloses the lensed source images. The pixelization grid for the free-form model of lensing potential perturbation is thus defined within this annular region, using a regular grid with a pixel size $N_{\delta\psi}$ times larger than that of the native lensing image [^4] This functionality is supported by the RegularDpsiMesh module in our code, where the factor parameter corresponds to $N_{\delta\psi}$., where $N_{\delta\psi}=1,2,\ldots$. The default $N_{\delta\psi}$ value used in this work is 2, which strikes a balance between computational efficiency and the accuracy required to recover the small mass perturbations induced by subhalos.
The Matérn regularisation (Stein, 1999; Rasmussen and Williams, 2006; Matern, 2013) is employed to control the smoothness of the pixelized lensing potential perturbation. For any two pixels of the lensing potential perturbation, denoted by their coordinates $(x_{m},y_{m})$ and $(x_{n},y_{n})$, the correlation between them can be described by
$C_{ij}(\rho,\nu)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\sqrt{2\nu}\frac{d_{ij}}{\rho}\right)^{\nu}K_{\nu}\left(\sqrt{2\nu}\frac{d_{ij}}{\rho}\right),$ (19)
where $\Gamma$ is the gamma function, and $K_{\nu}$ is the modified Bessel function of the second kind. $d_{ij}=\sqrt{(x_{m}-x_{n})^{2}+(y_{m}-y_{n})^{2}}$ represents the distance between these two pixels. The parameter $\rho$ defines the characteristic smoothing scale, while $\nu$ governs the order of smoothing, ensuring that the pixelized model reconstruction is $\left\lceil\nu\right\rceil-1$ times differentiable. The Exponential and Gaussian kernels used in Vernardos and Koopmans (2022) are special cases of the Matérn kernel, with $\nu=0.5$ and $\nu=+\infty$, respectively. The Hessian ($\mathbf{R}_{\delta\psi}$) of the regularisation function, using the Matérn scheme, is defined as follows:
@@ 3 Overview of Strong Lensing Data
Gaussian Random Field Normalization ($\log[A_{\mathrm{GRF}}]$) 6.0-6.0 $\left(-4.35^{0.03}_{-0.03}\right)$
@@ 3.1 Mock Lensing Data
We generate a set of mock lensing images that incorporate various perturbation signals in the lensing potential to evaluate the effectiveness of the potential correction code developed in this study. To control for variables, we ensure that the mass of the main lens galaxy and the source light remain consistent across all mock datasets, allowing only the perturbation signals in the lensing potential to vary. The mass distribution of the main lens galaxy is described by the elliptical power-law model (Tessore and Metcalf, 2015), whose dimensionless surface mass density (i.e. convergence) profile is expressed as
@@
Here, $1<\gamma<3$ represents the slope of the deprojected volume density, $\theta_{\mathrm{E}}$ denotes the Einstein radius, and $r$ signifies the elliptical radial distance from the centre of the lens galaxy. The elliptical radial distance is defined as $r(x,y)=\sqrt{x^{2}q+\frac{y^{2}}{q}}$, where $q$ is the minor-to-major axis ratio, and $\phi$ is the position angle measured counterclockwise from the positive x-axis. The parameters $q$ and $\phi$ are reparametrized into two ellipticity components, $\epsilon_{1}$ and $\epsilon_{2}$, defined as
@@
Four popular types of lensing potential perturbation signals, frequently used to evaluate the performance of free-form models (e.g. Vernardos and Koopmans, 2022; Galan et al., 2022; Biggio et al., 2023), are employed to construct mock datasets and test our potential correction code. These include: a spherical Navarro-Frenk-White (NFW) profile representing subhalos within the lens galaxy (Navarro et al., 1996), an external shear field caused by neighbouring galaxies or the cosmic large-scale structure (Keeton et al., 1997), a multipole perturbation induced by galaxy interactions such as mergers (Hao et al., 2006), and a Gaussian Random Field (GRF), which is expected to capture the collective lensing effects of numerous line-of-sight halos (Chatterjee and Koopmans, 2018; Bayer et al., 2023). The NFW model has a volume mass density profile of
@@
where $a_{m}$ represents the normalization factor, and $m$ and $\phi_{m}$ denote the multipole order and orientation, respectively (c.f. Congdon and Keeton, 2005; Taylor et al., 2017; Etherington et al., 2024; Paugnat and Gilman, 2025). The characteristics of a GRF are defined by its isotropic power spectrum, which follows a power-law relation:
@@ 3.2 Lensing Data for SLACS0946+1006
SLACS0946+1006 is a compound lens system with a rare triple source-plane configuration, characterized by a lens redshift of $z_{l}=0.222$ and source redshifts of $z_{s1}=0.609$, $z_{s2}=2.035$, and $z_{s3}=5.975$ (Smith and Collett, 2021). This system was initially pre-selected as a lens candidate from early-type galaxies in the SDSS survey, which exhibited abnormal emission lines potentially originating from high-redshift background sources (Bolton et al., 2006). The ACS camera on board the HST was subsequently employed to confirm the strong lensing nature of this system. The SLACS0946+1006 data analysed in this work were obtained using the ACS F814W band and then drizzled to $0.05\arcsec\text{pixel}^{-1}$, following the reduction procedures described in Bolton et al. (2008). The PSF was determined using the TinyTim software. For simplicity, we masked the outer Einstein ring, using only the innermost Einstein ring for potential correction. The lens light was pre-subtracted using the model results from He et al. (2024), which employs the Multipole Gaussian Expansion (MGE) model to represent the lens light; combined with a pixelized source model, the lens and source light are fit simultaneously to account for the covariance between the two. Subtracting the lens light using MGE typically outperforms multiple Sersic (Sérsic, 1963) fits in terms of image residuals. Figure 2 shows the lens-light-subtracted image of SLACS0946+1006.
@@ 4.1 Inverting Only the Potential Perturbation
Figure 3: Reconstruction of the lensing potential perturbation induced by a $10^{10}M_{\odot}$ NFW subhalo using the Matérn regularisation scheme, assuming that the true source and the image residuals induced by the perturber are perfectly known. Top-left: Image residual induced by the $10^{10}M_{\odot}$ NFW subhalo. Top-middle: Image residual reconstructed using the potential correction model. Top-right: Difference between the top-left and top-middle panels, normalised by the noise map. Bottom-left: Lensing potential map of the input perturber. Bottom-middle: Lensing potential map of the model perturber, reconstructed using the potential correction model. Bottom-right: Difference between the bottom-left and bottom-middle panels. The ‘plus’ symbol in the top-left and top-middle panels marks the position of the input subhalo.
@@
Figure 5: Comparison of reconstructed lensing potential perturbations induced by an NFW subhalo (left column), external shear (middle-left column), $m_{4}$ multipole (middle-right column), and a Gaussian random field (GRF) (right column) with the ground truth. The first row presents the input lensing potential maps of perturbers. The second row shows the perturbative lensing potential maps derived from the potential correction model. The third row presents a quantitative analysis: for the NFW subhalo, external shear, and $m_{4}$ multipole, parametric models are fitted to the lensing potential map in the second row, with point estimates reported in Table 3; for the GRF, the power spectra of the input (blue lines) and reconstructed (red lines) lensing potential maps are compared. Solid lines indicate power spectra calculated within the unmasked modelling region, while dashed lines represent power-law model fits.
@@
Parameter Names $\delta\boldsymbol{\psi}$-Only (§ 2.2) $(\delta\boldsymbol{\psi},\boldsymbol{s})^{\mathrm{Truth}}$ (§ 4.2.1) $(\delta\boldsymbol{\psi},\boldsymbol{s})^{\mathrm{Macro}}$ (§ 4.2.2) Ground Truths
@@
Mass ($\log_{10}[M_{200}/\mathrm{M}_{\odot}]$) $9665^{0.7155}_{-0.6685} 9.9$$9151^{0.7237}_{-0.6488}$ $9.5192^{0.8085}_{-0.3837}$ 10.0
@@
Table 3: Quantitative comparison of reconstructed and input lensing potential perturbations. For subhalo, shear, and multipole models, parametric models are fitted to the reconstructed perturbations to assess agreement with the input. For the Gaussian random field, the power spectrum of the perturbations is compared. The labels $\delta\boldsymbol{\psi}$-Only, $(\delta\boldsymbol{\psi},\boldsymbol{s})^{\mathrm{Truth}}$, and $(\delta\boldsymbol{\psi},\boldsymbol{s})^{\mathrm{Macro}}$ denote inversions of only the lensing potential perturbations (§ 2.2), inversions of both source and lensing potential perturbations starting from the exact true main lens mass and source light distributions (§ 4.2.1), and inversions of both source and lensing potential perturbations starting from the macro model fit (§ 4.2.2), respectively. Superscripts and subscripts represent the 68% credible interval.
@@
To better visualise the structure of a localised subhalo, the convergence map is more informative than the lensing potential map. The top-left and top-right panels of Figure 4 compare the convergence maps of the input perturber and the one derived by the potential correction model, respectively. Our potential correction method with Matérn regularisation successfully reconstructs the mass distribution of the input NFW subhalo qualitatively. In contrast, using Exponential or Gaussian regularisation, as proposed in Vernardos and Koopmans (2022), leads to under- or over-smoothed solutions that fail to recover the input NFW subhalo, as shown in the bottom-left and bottom-right panels of Figure 4. This demonstrates the superiority of Matern regularisation in reconstructing localised subhalo perturbations.
To enable a more quantitative assessment of the potential correction method, we fit a parametric model to the perturbative lensing potential map derived by the potential correction method for mock lenses with an NFW subhalo [^5] The NFW subhalo is a localised perturber. For this parametric model fitting, we only use the potential correction results in the central region (indicated by the black circle in Figure 4), as the outskirts are likely degraded by noise., external shear, and an $m_{4}$ multipole perturber, and evaluate whether the fitting results agree with the true input values. For the Gaussian random field, we compare the power spectrum of the model-derived lensing potential perturbations with the true input values. As shown in Figure 5, the lensing potential perturbations derived by the potential correction method (second row) are consistent with the true input values (first row). A quantitative diagnostic (bottom row), conducted either via parametric model fitting or power spectrum comparison, demonstrates that our potential correction method can accurately recover the input perturbations. The specific parametric values from this diagnostic are summarised [^6] It should be noted that, although the power spectrum properties of the input Gaussian random field are recovered reasonably well using the potential correction method, the model and the input true values do not perfectly align within the statistical uncertainty, a phenomenon also observed in Vernardos and Koopmans (2022). in Table 3.